Can someone please help me solve this algebra problem. I set up the problem the best way, I knew how.
1/4 =3- 2x-1/x+2
This problem is suppose to be set up as a fraction 1/4 and 2x-1/x+2

Answer by stanbon(75887)

(1/4) =3- [(2x-1)/(x+2)]
-------
Multiply thru by 4(x+2) to get:
x+2 = 12(x+2) - (2x-1)
-----
x+2 = 12x + 24 -3x + 1
-----
x+2 = 9x + 25
8x = -23
x = -23/8
===============
Cheers,
Stan H.
=====================